What is the weight/volume percent concentration of sodium chloride in seawater if one liter contains 35 g of NaCl?
1
Expert's answer
2020-09-14T08:23:42-0400

%W/V={Wt of solute in(gram)/Volume

of solution in (ml)}×100

Given,1 liter=1000ml Vol. of solution

35 gram = weight of solute

We get,

%W/V=(35/1000)×100=3.5 ans